# sample.bib

### v2

@@ Abstract
Methods: A set of 100 multiple choice radiation oncology physics questions, previously created by a well-experienced physicist, was used for this study. The answer options of the questions were randomly shuffled to create ”new” exam sets. Five LLMs (OpenAI o1-preview, GPT-4o, LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet) with the versions released before September 30, 2024, were queried using these new exam sets. To evaluate their deductive reasoning capabilities, the correct answers in the questions were replaced with ”None of the above.” Then, the explaining-first and step-by-step instruction prompts were used to test if this strategy improved their reasoning capabilities. The performance of the LLMs was compared with the answers from medical physicists.
Results: All models demonstrated expert-level performance on these questions, with o1-preview even surpassing medical physicists with a majority vote. When replacing the correct answers with ’None of the above’, all models exhibited a considerable decline in performance, suggesting room for improvement. The explaining-first and step-by-step instruction prompts helped enhance the reasoning capabilities of the LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet models.
Conclusion: These recently released LLMs demonstrated expert-level performance in answering radiation oncology physics questions, exhibiting great potential to assist in radiation oncology physics training and education.
@@ 1 Introduction
Large language models (LLMs) have advanced rapidly. On the one hand, the size of the data used for pre-training and the number of model parameters have grown significantly [gururangan-etal-2020-dont, kaplan2020scalinglawsneurallanguage]; on the other hand, the fine-tuning methods and prompt engineering strategies have advanced substantially [ouyang2022traininglanguagemodelsfollow, hu2021loralowrankadaptationlarge]. Furthermore, agents and Retrieval-Augmented Generation (RAG) systems built on LLMs have seen considerable progress [xi2023risepotentiallargelanguage, lewis2021retrievalaugmentedgenerationknowledgeintensivenlp].The notable recent developments include OpenAI o1-preview [o1_webpage], GPT-4o [openai2024gpt4ocard], LLaMA 3.1 (405) [grattafiori2024llama3herdmodels], Gemini 1.5 Pro [geminiteam2024gemini15unlockingmultimodal], and Claude 3.5 Sonnet [claude_webpage], each demonstrating remarkable improvements in language comprehension, reasoning, and domain-specific expertise.
@@
The rapid evolution of LLMs also renders prior performance evaluations outdated. As some LLMs cease providing services, new models are introduced, and existing versions are updated, studies published before may no longer accurately reflect the current state of LLM capabilities. A fresh evaluation is needed to address the dynamic landscape due to the rapid LLM development.
@@
Although,large language models (LLMs) have shown great potential for applications in healthcare recently [liuzhengliang2023, Liu2024a, LIU2023100045, Liu2022SurveyON, Rezayi2022, zhangkai2024, zhang2024, Dai2023ChatAugLC, HOLMES2024e515, XIAO2024102204, Wu2023ExploringTT, LIAO2024128576, li2024, dai2023samaug], their direct use in specialized fields such as radiation oncology presents challenges and necessitates thorough evaluation and validation. Radiation oncology is distinct from other clinical practices due to two salient features: a high degree of complexity and stringent precision requirements [dawson2007advances, xing2006overview]. This is particularly true for advanced radiation therapy modalities like intensity-modulated X-ray-based radiation therapy and intensity-modulated proton therapy [quan2013preliminary, cao2012uncertainty, zaghian2017comparison, liu2018impact, an2017robust, liu2020robust, shan2020intensity]. Therefore, assessing the performance of LLMs in addressing questions related to radiation oncology physics is crucial for their successful integration into this field. Such evaluations not only show us how effectively they process and understand radiation oncology physics, but also help us understand their limitations. This provides us with guidance in this field, especially for their potential utilization in training and education in radiation oncology physics. Built on our prior work [10.3389/fonc.2023.1219326], we provided an updated study with refined methods, offering insights into the performance of the recently released LLMs in radiation oncology physics.
@@
We utilized the 100-question radiation oncology physics exam we developed based on the American Board of Radiology exam style [liu2024radiationoncologynlpdatabase], and randomly shuffled the answer options to create ”new” exam sets. We then queried the LLMs with these new exam sets and checked their capability to answer questions accurately. We also evaluated their deductive reasoning capabilities and tested whether the explaining-first and step-by-step instruction prompts would improve their performance in answering radiation oncology physics questions.
@@ 2 Methods
The 100 multiple-choice exam questions on radiation oncology physics were created by our experienced medical physicist, following the official study guide of American Board of Radiology. That exam includes 12 questions on basic physics, 10 questions on radiation measurements, 20 questions on treatment planning, 17 questions on imaging modalities and applications in radiotherapy, 13 questions on brachytherapy, 16 questions on advanced treatment planning and special procedures, and 12 questions on safety, quality assurance (QA), and radiation protection. 17 out of the 100 questions are math-based and require numerical calculation.
@@ 2.1 Randomly shuffling the answer options
A Python code was developed to randomly shuffle the answer options for the 100 multiple-choice questions five times. For each shuffle, we obtained a ”new” 100-question multiple-choice exam set. We then queried all the LLMs five times. For each time (Trial 1 - Trial 5), we queried all the LLMs with a ”new” exam set generated by randomly shuffling. Each question of every exam set was queried individually. We checked the distribution of the correct answers’ locations for the five exam sets and confirmed that the distribution of the correct options is fairly random among A, B, C, D, or E (only 2 questions offered option E).
@@
This allowed the LLMs to reason and answer freely. Table 1 illustrates an example of the trials and how we queried the LLMs with the questions. For the responses generated by the LLMs, we utilized the LLaMA 3.1 (405B) model hosted locally to further extract the chosen answer options (letter A, B, C, D, or E) from responses. We then conducted manual verification and correction of the extracted options and obtained the final answer option sheet for all LLMs. The accuracy of each LLM was reported as the mean score across the five trials, and the measurement uncertainty was reported as the standard deviation of the five trials, as shown in the following equations:
@@
2.2 Evaluating deductive reasoning capability
Deductive reasoning capability refers to the cognitive process of logically analyzing information to draw specific conclusions from general premises. The multiple-choice question with the answer option ”None of the above” can effectively evaluate the test-taker’s deductive reasoning capability, as it involves evaluating each option based on the information provided and ruling out incorrect choices contradicting known facts or logical outcomes to reach the correct answer. We therefore had replaced the correct option in the exam with ”None of the above.” Since transformer-based LLMs predict the next word based on prior contexts, changing the correct option to ”None of the above” removes a straightforward cue that might guide the model toward a known or patterned solution, thus forcing the LLMs to rely more on reasoning about the specific question and its options, rather than using surface-level lexical or statistical patterns that it may have learned.
@@ 2.2.1 Replacing the correct option with ”None of the above”
A Python code was developed to replace the correct option with ”None of the above” for all questions in the five new exam sets derived by random shuffling. For each trial (Trial 1 - Trial 5), we queried the LLMs with a set of exams in which both the correct option was ”None of the above,” and its location was randomly shuffled. We used the same prompt as in Sec. 2.1 across all queries, and each question in an exam set was queried individually. This setup challenges the LLMs to avoid pattern-based answering and not rely on any single choice, but to process each answer option by reasoning.
Similar to before, we utilized the previously described processes for answer-option extraction, manual verification, and correction, as outlined in Sec. 2.1. The performance accuracy and uncertainty of each LLM were reported as the average score and standard deviation across all five trials. Due to this setup, these exams were not used to test humans, as this pattern can be easily recognized by human test-takers [^2] Since each question was queried through the API individually, it is assumed that LLMs would not notice this pattern.
2.2.2 Explaining first and step-by-step instruction
To further check if explicitly asking the LLMs to explain first and then develop answers step-by-step would improve their deductive reasoning capability [wei2023chainofthoughtpromptingelicitsreasoning], we engineered the following prompt and queried the LLMs again with it:
@@ 2.2.2 Explaining first and step-by-step instruction
We used the five exam sets and conducted the querying process both as described in Sec. 2.2.1. All five LLMs were evaluated using this prompting strategy. Accuracy and uncertainty were reported as before. The results from this strategy were compared with the test results from original prompts, where no explanation or step-by-step answering was required, as described in Sec. 2.2.1.
@@ 3.1 Results of exam sets with randomly shuffled options
The evaluation results of the exam sets with randomly shuffled options are presented in Fig. 1, where the height of each bar represents the mean test score, and the error bars indicate the standard deviations across five trials. All five LLMs exhibited strong performance, achieving mean test scores above 80%, which suggests their performance on these exams is comparable to that of human experts. When compared to the majority vote results from the medical physics group, the OpenAI o1-preview model outperformed the medical physicists with a majority vote. For math-based questions, both the o1-preview and GPT-4o models surpassed the medical physicists with a majority vote.
@@
Figure 1: Evaluation results using five exam sets, where the answer options of the questions were randomly shuffled. Figure (a) illustrates the evaluation results for all questions and math-based questions, while Figure (b) presents the evaluation results broken down by different topics.
@@
The raw counts of incorrect answers by the LLMs are shown in Fig. 2, where each color represents the incorrect answers by an LLM across trials. As observed, each LLM exhibited variability in answering questions across trials. Notably, the models also showed similarities in incorrectly answering certain questions. We analyzed the questions that all five models had ever answered incorrectly across all five trials, i.e., question numbers: 14, 27, 33, 42, 67, 87, 95, 96. Interestingly, only one of these questions was math-based, while the remaining seven were closely related to clinical medical physics knowledge, such as American Association of Physicists in Medicine (AAPM) Task Group (TG) reports and clinical experience. This observation suggests that current LLMs may still struggle with answering clinically-focused radiation oncology physics questions. For example, question number 42 does not involve any calculations but instead focuses primarily on clinical hands-on experience.
@@
Figure 2: Distribution of incorrect answer counts of every LLM across five exam sets, where answer options were randomly shuffled. Questions that were correctly answered by LLMs in all five trials are not displayed in this figure.
3.2 Results of LLMs’ deductive reasoning capability
@@
Fig. 3 shows the results of the deductive reasoning capability tests, where the correct answer options were replaced with ”None of the above” in all questions. Overall, all LLMs performed much more poorly compared to the results in Sec. 3.1. Given that transformer-based LLMs [vaswani2023attentionneed] were designed to predict the next word in a sequence, replacing the correct answers with ”None of the above” would likely disrupt their pattern recognition capabilities, thereby reducing their overall scores performed on the exam sets. Nonetheless, the OpenAI o1-preview and GPT-4o still outperformed the others, especially on math-based questions, indicating strong reasoning capabilities of these two models.
@@
Figure 3: Deductive reasoning capability evaluation results for every LLM. The correct answer options were replaced with ”None of the above” in all questions. Figure (a) illustrates the evaluation results for all questions and math-based questions, while Figure (b) presents the evaluation results broken down by different topics.
@@
Fig. 4 compares the performance of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet models with the original simple prompts and with the explaining-first, step-by-step instruction prompts. Overall, all three models demonstrated improved reasoning capabilities with the latter prompting strategy. Notably, Gemini 1.5 Pro showed significant gains on math-based questions, increasing its score from 24% to 68%. The o1-preview and GPT-4o improved by only about 1% overall, too small to be represented in this figure.
@@
Figure 4: Improvements in reasoning capabilities observed using the explaining-first and step-by-step instruction prompts for the LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. Figure (a) illustrates the improvements for all questions and math-based questions, while Figure (b) presents the improvements broken down by different topics. (Note: The o1-preview and GPT-4o showed only about 1% overall improvement, too small to be represented in this figure.)
@@ 4 Discussion
Over the past two years, our studies have observed a notable improvement in the performance of state-of-the-art LLMs on this highly specialized task - answering radiation oncology physics questions, as shown in Fig. 5. Earlier versions of ChatGPT, like GPT-3.5 in late 2022 [chatgpt_webpage], scored around 54%, showing clear gaps in domain-specific knowledge. With the introduction of GPT-4 in early 2023 [openai2024gpt4technicalreport], performance jumped to around 76%, reflecting improvements in accuracy and understanding. Subsequent releases of the GPT-4o model and more recently the o1-preview (both in 2024), pushed scores even higher to 90% and 94% respectively, indicating improved domain knowledge in radiation oncology physics. This steady improvement can be attributed to more extensive domain pre-training, refined architectural updates [brown2020languagemodelsfewshotlearners], and enhanced fine-tuning techniques, all of which have led to stronger reasoning skills, improved understanding, and better alignment with expert-level knowledge. The evolution of these models over the last two years underscores the rapid growth of LLMs, suggesting their potential as useful tools in areas such as radiation oncology physics training and education.

### v3

@@ Abstract
Methods: A set of 100 multiple-choice radiation oncology physics questions, previously created by a well-experienced physicist, was used for this study. The answer options of the questions were randomly shuffled to create ”new” exam sets. Five LLMs – OpenAI o1-preview, GPT-4o, LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet – with the versions released before September 30, 2024, were queried using these new exam sets. To evaluate their deductive reasoning ability, the correct answer options in the questions were replaced with ”None of the above.” Then, the explain-first and step-by-step instruction prompts were used to test if this strategy improved their reasoning ability. The performance of the LLMs was compared with the answers from medical physicists.
Results: All models demonstrated expert-level performance on these questions, with o1-preview even surpassing medical physicists with a majority vote. When replacing the correct answer options with ’None of the above’, all models exhibited a considerable decline in performance, suggesting room for improvement. The explain-first and step-by-step instruction prompts helped enhance the reasoning ability of the LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet models.
Conclusion: These recently released LLMs demonstrated expert-level performance in answering radiation oncology physics questions, exhibiting great potential to assist in radiation oncology physics education and training.
@@ 1 Introduction
Large language models (LLMs) have advanced rapidly. On the one hand, the size of the data used for pre-training and the number of model parameters have grown significantly [gururangan-etal-2020-dont, kaplan2020scalinglawsneurallanguage]; on the other hand, the fine-tuning methods and prompt engineering strategies have advanced substantially [ouyang2022traininglanguagemodelsfollow, hu2021loralowrankadaptationlarge]. Furthermore, agents and Retrieval-Augmented Generation (RAG) systems built on LLMs have seen considerable progress [xi2023risepotentiallargelanguage, lewis2021retrievalaugmentedgenerationknowledgeintensivenlp]. The notable recent developments include OpenAI o1-preview [o1_webpage], GPT-4o [openai2024gpt4ocard], LLaMA 3.1 (405B parameters) [grattafiori2024llama3herdmodels], Gemini 1.5 Pro [geminiteam2024gemini15unlockingmultimodal], and Claude 3.5 Sonnet [claude_webpage], each demonstrating remarkable improvements in language comprehension, reasoning, and domain-specific expertise.
@@
The rapid evolution of LLMs also renders prior performance evaluations outdated. As some LLMs cease providing services, new models are introduced, and existing versions are updated, studies published before may no longer accurately reflect the current state of LLM capabilities. A fresh evaluation is needed to address the dynamic landscape of LLM advancements.
@@
In healthcare, LLMs have been explored for numerous potential applications [Liu2024a, LIU2023100045, zhangkai2024, HOLMES2024e515, li2024, liu2024automatedradiotherapytreatmentplanning]. For their direct use in radiation oncology, unique challenges related to evaluation and validation are posed – due to the high level of complexity and stringent precision required [dawson2007advances, xing2006overview]. This is particularly true for advanced radiation therapy modalities like intensity-modulated X-ray-based radiation therapy and intensity-modulated proton therapy [quan2013preliminary, cao2012uncertainty, zaghian2017comparison, liu2018impact, an2017robust, liu2020robust, shan2020intensity]. Therefore, assessing the performance of LLMs in addressing questions related to radiation oncology physics is crucial. Such evaluations not only tell us how efficiently they process and reason about radiation oncology physics but also help us understand their limitations. This provides us with guidance in this field, especially for their potential utilization in education and training. Built on our prior work [10.3389/fonc.2023.1219326], we provided an updated study with refined methods, offering insights into the performance of the recently released LLMs in radiation oncology physics.
@@
We utilized the 100-question radiation oncology physics exam we developed based on the American Board of Radiology exam style [liu2024radiationoncologynlpdatabase], and randomly shuffled the answer options to create ”new” exam sets. We then queried the LLMs with these new exam sets and checked their ability to answer questions accurately. We also evaluated their deductive reasoning ability and tested whether the explain-first and step-by-step instruction prompts would improve their performance in reasoning tasks.
@@ 2 Methods
The 100-question multiple-choice examination on radiation oncology physics was created by our experienced medical physicist, following the official study guide of the American Board of Radiology. That exam includes 12 questions on basic physics, 10 questions on radiation measurements, 20 questions on treatment planning, 17 questions on imaging modalities and applications in radiotherapy, 13 questions on brachytherapy, 16 questions on advanced treatment planning and special procedures, and 12 questions on safety, quality assurance (QA), and radiation protection. 17 out of the 100 questions are math-based and require numerical calculation.
@@ 2.1 Randomly shuffling the answer options
Since it was difficult to know whether any LLM had been pre-trained using our previously published 100-question multiple-choice exam, we wrote Python code to randomly shuffle the answer options for the 100 multiple-choice questions five times. For each shuffle, we obtained a ”new” 100-question multiple-choice exam set. We then queried all the LLMs five times (Trial 1 - Trial 5), each with a new exam set. Each question of the new exam set was queried individually. We checked the distribution of the correct answers’ locations for the five new exams where the options were shuffled and confirmed that the distribution of the correct options is fairly random among A, B, C, D, or E (only 2 questions offered option E), as shown in the supplementary material.
@@
This allowed the LLMs to reason and answer freely. Table 1 illustrates an example of the trials and how we queried the LLMs with the questions. For the responses generated by the LLMs, we utilized the LLaMA 3.1 (405B) model hosted locally to further extract the chosen answer options (letters A, B, C, D, or E) from responses. We then conducted manual verification and correction of the extracted options and obtained the final answer option sheet for all LLMs. The accuracy of each LLM was reported as the mean score across the five trials, and the measurement uncertainty was reported as the standard deviation of the five trials, as shown in the following equations:
@@
2.2 Evaluating deductive reasoning ability
Deductive reasoning ability refers to the cognitive process of logically analyzing information to draw specific conclusions from general premises. The multiple-choice question with the answer option ”None of the above” can effectively evaluate the test-taker’s deductive reasoning ability, as it involves evaluating each option based on the information provided and ruling out incorrect choices contradicting known facts or logical outcomes to reach the correct answer. We therefore replaced the correct option in the exam with ”None of the above.” Since transformer-based LLMs predict the next word based on prior contexts, changing the correct option to ”None of the above” removes a straightforward cue that might guide the model toward a known or patterned solution, thus forcing the LLMs to rely more on reasoning about the specific question and its options, rather than using surface-level lexical or statistical patterns that it may have learned.
@@ 2.2.1 Replacing the correct option with ”None of the above”
We developed Python code to replace the correct option with ”None of the above” for all questions in the five new exam sets derived by random shuffling. For each trial (Trial 1 - Trial 5), we queried the LLMs with a set of exams in which both the correct option was ”None of the above,” and its location was randomly shuffled. We used the same prompt as in Sec. 2.1 across all queries, and each question in an exam set was queried individually. This setup challenges the LLMs to avoid pattern-based answering and not rely on any single choice, but to process each answer option by reasoning.
As before, we utilized the previously described processes for answer-option extraction, manual verification, and correction, as outlined in Sec. 2.1. The performance accuracy and uncertainty of each LLM were reported as the average score and standard deviation across all five trials. Due to this setup, these exams were not used to test humans, as this pattern can be easily recognized by human test-takers [^2] Since each question was queried through the API individually, it is assumed that LLMs would not notice this pattern.
2.2.2 Explain-first and step-by-step instruction
To further check if explicitly asking the LLMs to explain first and then develop answers step-by-step would improve their deductive reasoning ability [wei2023chainofthoughtpromptingelicitsreasoning], we engineered the following prompt and queried the LLMs again with it:
@@ 2.2.2 Explain-first and step-by-step instruction
We used the five exam sets and conducted the querying process both as described in Sec. 2.2.1. All five LLMs were evaluated using this prompting strategy. Accuracy and uncertainty were reported. The results from this strategy were compared with the test results from original prompts, where no explanation or step-by-step answering was required, as described in Sec. 2.2.1.
@@ 3.1 Results of exam sets with randomly shuffled options
The evaluation results of the exam sets with options randomly shuffled are presented in Fig. 1, where the height of each bar represents the mean test score, and the error bars indicate the standard deviations across five trials. All five LLMs exhibited strong performance, achieving mean test scores above 80%, which suggests their performance on these exams is comparable to that of human experts. When compared to the majority vote results from the medical physics group, the OpenAI o1-preview model outperformed the medical physicists with a majority vote. For math-based questions, both the o1-preview and GPT-4o models surpassed the medical physicists with a majority vote.
@@
Figure 1: Evaluation results using five exam sets where the answer options of the questions were randomly shuffled. Figure (a) illustrates the evaluation results for all questions and math-based questions, while Figure (b) presents the evaluation results broken down by different topics.
@@
The raw counts of incorrect responses by the LLMs are shown in Fig. 2, where each color represents the incorrect answers by an LLM across trials. As observed, each LLM exhibited variability in answering questions across trials. Notably, the models also showed similarities in incorrectly answering certain questions. We analyzed the questions that all five models had ever answered incorrectly across all five trials – question numbers: 14, 27, 33, 42, 67, 87, 95, 96. Interestingly, only one of these questions was math-based, while the remaining seven were closely related to clinical medical physics knowledge, such as American Association of Physicists in Medicine (AAPM) Task Group (TG) reports and clinical experience. This observation suggests that current LLMs may still struggle with answering clinically focused radiation oncology physics questions. For example, question number 42 does not involve any calculations but instead focuses primarily on clinical hands-on experience.
@@
Figure 2: Distribution of incorrect answer counts of each LLM across five exam sets where answer options were randomly shuffled. Questions that were correctly answered by LLMs in all five trials are not displayed in this figure.
3.2 Results of LLMs’ deductive reasoning ability
@@
Fig. 3 shows the results of the deductive reasoning ability tests, where the correct answer options were replaced with ”None of the above” in all questions. Overall, all LLMs performed much more poorly compared to the results in Sec. 3.1. Given that transformer-based LLMs [vaswani2023attentionneed] were designed to predict the next word in a sequence, replacing the correct answers with ”None of the above” would likely disrupt their pattern recognition abilities, thereby reducing their overall scores performed on the exam sets. Nonetheless, the OpenAI o1-preview and GPT-4o still outperformed the others, especially on math-based questions, indicating the strong reasoning ability of these two models.
@@
Figure 3: Deductive reasoning ability evaluation results for every LLM. The correct answer options were replaced with ”None of the above” in all questions. Figure (a) illustrates the evaluation results for all questions and math-based questions, while Figure (b) presents the evaluation results broken down by different topics.
@@
Fig. 4 compares the performance of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet models with the original simple prompts and with the explain-first, step-by-step instruction prompts. Overall, all three models demonstrated improved reasoning ability with the latter prompting strategy. Notably, Gemini 1.5 Pro showed significant gains on math-based questions, increasing its score from 24% to 68%. The o1-preview and GPT-4o improved by only about 1% overall, which is too small to be represented in this figure.
@@
Figure 4: Improvements in reasoning abilities observed using the explain-first and step-by-step instruction prompts for the LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. Figure (a) illustrates the improvements for all questions and math-based questions, while Figure (b) presents the improvements broken down by different topics. (Note: The o1-preview and GPT-4o showed only about 1% overall improvement, too small to be represented in this figure.)
@@ 4 Discussion
Over the past two years, our studies have observed a notable improvement in the performance of state-of-the-art LLMs on this highly specialized task – answering radiation oncology physics questions, as shown in Fig. 5. Early versions of ChatGPT, like GPT-3.5 in late 2022 [chatgpt_webpage], scored around 54%, showing clear gaps in domain-specific knowledge. With the introduction of GPT-4 in early 2023 [openai2024gpt4technicalreport], performance jumped to around 76%, reflecting improvements in accuracy and understanding. Subsequent releases of the GPT-4o model and more recently the o1-preview (both in 2024), pushed scores even higher to 90% and 94% respectively, indicating increasing capabilities in radiation oncology physics. This steady improvement can be attributed to more extensive domain pre-training, refined architectural updates [brown2020languagemodelsfewshotlearners], and enhanced fine-tuning techniques, all of which have led to stronger reasoning skills, improved understanding, and better alignment with expert-level knowledge. The evolution of these models over the last two years underscores the rapid growth of LLMs, suggesting their potential as useful tools in areas such as radiation oncology physics education and training.